OpenStudy (anonymous):

cos(-11pi/4) sin(-11pi/4)

OpenStudy (anonymous):

\[\cos(a+k2\pi)=\cos(a) \]\[\sin(a+k2\pi)=\sin(a) \] for all integers k

OpenStudy (anonymous):

So basically you can factor out multiples of 2 pi

OpenStudy (anonymous):

yeah, i tried that i thought the exact value of cos what square root 3 and sin was 1/2

OpenStudy (anonymous):

are you asking\[\cos(-11\frac{ \pi }{ 4 })\]or\[\cos(\frac{ -11\pi }{ 4 })\]

OpenStudy (anonymous):

-11pi/4

OpenStudy (anonymous):

That doesn't clarify. Is it (-11 pi)/4 or -11 (pi/4)

OpenStudy (anonymous):

(-11 pi)/4

OpenStudy (anonymous):

So you subtract 2pi from that and what do you get?

OpenStudy (anonymous):

Sorry add 2pi

OpenStudy (anonymous):

13pi/4

OpenStudy (anonymous):

\[\frac{ -11\pi }{ 4 }+2\pi=\frac{ -11\pi }{ 4 }+\frac{ 8 \pi }{ 4 }=\frac{ -3 \pi }{ 4 }\]

OpenStudy (anonymous):

These are some values to remember:\[\cos (\pi/4)=\sin (\pi/4)=\frac{ \sqrt{2} }{ 2 }\] so that is for a 45 degree angle in quadrant I

OpenStudy (anonymous):

Which quadrant does (-3/4)pi lie in?